Tom is 4 years older than his brother, Ray. The product of their ages is 45. Find the age of each brother.
Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

Find a pair of factors of 45 that differ by 4.
====================
Or: x = Ray's age
---
x(x+4) = 45
x^2 + 4x - 45 = 0
At this point, find a pair of factors of 45 that differ by 4.

Answer by ewatrrr(24785)



Hi
Let x be Ray's age
Question States:
 x(x+4) = 45
 x^2 + 4x -45 = 0
 (x-5)(x+9) = 0    Tossing out the negative value for age
   x = 5, Ray's age
and
   Tom is 9
 
  5*9 = 45 checks

Wish You the Best in your Studies.